what is the formula for a triangular prism? thank you 4 ur time & answer

p.s.: wat is the formula 4 the volume of a trapezodial prism? and the 2 steps to find the volume f any prism or cylinder???
thank u soooo much

Answer by checkley77(12569)

volume=1/2*b*h
-----------------
volume=base area*Height
volume=1/2h(b+B)H: where h=height of trapezoid, b&B are the 2 bases & H is the height of the trapezoid.
---------------------------------
volume=bh
volume=(pi*r^2)h